OpenStudy (anonymous):

Find cot θ if csc θ = square root of seventeen divided by four and tan θ > 0

OpenStudy (jdoe0001):

\(\bf csc(\theta) = \cfrac{\sqrt{17}}{4}\quad ?\)

OpenStudy (anonymous):

im sorry what?

OpenStudy (anonymous):

yes it has that then after that it says "and tan θ > 0"

OpenStudy (jdoe0001):

ok

OpenStudy (anonymous):

you know how to do it?

OpenStudy (jdoe0001):

well notice that \(\bf csc(\theta) = \cfrac{\sqrt{17}}{4} \implies \cfrac{\textit{hypotenuse}}{\textit{opposite side}} \implies \cfrac{c}{b}\\\quad \\ \textit{now get the "a" or adjacent side, keeping in mind that }\\ c^2 = a^2 + b^2 \implies \sqrt{c^2-b^2} = a\)

OpenStudy (jdoe0001):

\(\bf csc(\theta) = \cfrac{\sqrt{17}}{4} \implies \cfrac{\textit{hypotenuse}}{\textit{opposite side}} \implies \cfrac{c}{b}\\\quad \\ c = \sqrt{17} \qquad \qquad b = 4\)

OpenStudy (anonymous):

ok I see that what do I do to start solving it

OpenStudy (anonymous):

thank you btw

OpenStudy (jdoe0001):

well, you see, you're asked for the cotangent function the cotangent is \(\bf cot(\theta) = \cfrac{adjacent}{opposite} \implies \cfrac{a}{b}\) so, we would need the "a" component, that is, the adjacent side, to get the cotangent function

OpenStudy (jdoe0001):

we know what "c" is, we know what "b" is, so, using the pythagorean theorem \(\bf c^2 = a^2 + b^2 \implies \sqrt{c^2-b^2} = a\)

OpenStudy (anonymous):

so its 1?

OpenStudy (jdoe0001):

hmmm, well, yes, is unit 1 now, we also know that the tangent of that angle is > 0, that is, is POSITIVE the tangent is only positive in 1st Quadrant and 3rd Quadrant now, also notice, the hypotenuse is always positive so if say \(\bf csc(\theta) = \cfrac{\sqrt{17}}{4} \implies \cfrac{\sqrt{17}}{+4}\) that means that the side "b" is positive, side "b" is the y-coordinate, so our "y" is positive now, where between 1st and 3rd Quadrant is the "y" positive? well in the 1st one so our angle is in the 1st Quadrant, that means the "a" side, that is our x-coordinate, is also positive on that quadrant, so our 1 from the pythagorean theorem, is +1 so, what do you think is our cotangent?

OpenStudy (anonymous):

ok im a little lost where ami looking to find the answer?

OpenStudy (jdoe0001):

hmmm... .well, check your trig identities for cotangent

OpenStudy (anonymous):

is it 1/4 the answer?

OpenStudy (jdoe0001):

\(\bf csc(\theta) = \cfrac{\sqrt{17}}{4} \implies \cfrac{\textit{hypotenuse}}{\textit{opposite side}} \implies \cfrac{c}{b}\\\quad \\ c = \sqrt{17} \qquad \qquad b = \color{red}{4}\qquad \qquad a = \color{blue}{1}\\\quad \\ cot(\theta) = \cfrac{adjacent}{opposite} \implies \cfrac{a}{b} \implies \cfrac{\color{blue}{1}}{\color{red}{4}}\)

OpenStudy (jdoe0001):

yeap

OpenStudy (anonymous):

sweet thank you :)

OpenStudy (jdoe0001):

yw